When 110j of heat is added to a gaseous system internal energy increases by 40 j the amnt of work done is
 
 

Grade:12th pass

1 Answers

Arun
4 years ago
D ar student

DQ = DU + DW

Þ DW = DQ - DU = 110 - 40 = 70 J

 

Hope it helps

 

Regards

Arun (askIITians forum expert)